# Graph-Informed Neural Networks for Sparse Grid-Based Discontinuity Detectors

### v1

@@ 3.1 Algorithm for Discontinuity Detection
Then, from (18) the theses hold. Indeed, for each $\Lambda_{\min}>0$, there exists a minimum $k\in\mathbb{N}$ such that $\Lambda/(2^{k+1})<\Lambda_{\min}$; therefore, there exists $h\in\mathbb{N}$, $h\leq k$, such that $\lambda^{(h)}<\Lambda_{\min}$ and the algorithm has a finite termination. Moreover, it holds
@@
$\mathrm{dist}(\boldsymbol{x}^{(h)},\delta)=\|{\boldsymbol{x}^{(h)}-\boldsymbol{x}^{(h)\,\delta}}\|\leq\frac{\lambda^{(h)}}{2}<\frac{\Lambda_{\min}}{2}\leq\frac{2\varepsilon}{2}\,.$
@@
A discontinuity detector based on a sparse grid $\mathcal{S}$ in $\mathbb{R}^{n}$ can be used also for detecting discontinuities of functions $g^{\prime}:\Omega^{\prime}\subseteq\mathbb{R}^{n^{\prime}}\rightarrow\mathbb{R}$, with $n^{\prime}<n$. Indeed, we can define a function $g:\Omega^{\prime}\times\mathbb{R}^{n-n^{\prime}}\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}$ as
@@
for each $\boldsymbol{x}\in\Omega^{\prime}\times\mathbb{R}^{n-n^{\prime}}$, and use the detector with respect to $g$. Moreover, for practical advantages, in this case Algorithm 1 can be modified to ignore any detected troubled point with at least one coordinate $x_{n^{\prime}+1},\ldots,x_{n}\neq 0$, assuming to start with sparse grids all centered in points with coordinates $x_{n^{\prime}+1}=\cdots=x_{n}=0$. The study of such a kind of generalization is postponed to future work.
@@ 4 NN-based Discontinuity Detectors
First, we will sketch the entire procedure in Section 4.1. Then, in the following subsections, we will provide details about the loss function and NN models. We refer the reader to Appendix A for further details related to the building of a deterministic approximation of an exact detector (Section A.1), the creation of synthetic data for the training (Section A.2), data preprocessing (Section A.3), and NN architecture archetypes (Section 4.3).
@@ 4.1 Building Procedure for NN-based Detectors
    Choose a space dimension $n$ and a sparse grid $\mathcal{S}$ in $\mathbb{R}^{n}$. We denote by $N$ the number of points in $\mathcal{S}$.
@@
    For each $g^{(q)}\in\mathcal{G}$, run Algorithm 1 with a detector $\Delta$ that is exact ($\Delta\equiv\Delta^{*}$) or it is a deterministic approximation of $\Delta^{*}$ (see Section A.1 for details); then, at each detection step, collect:
    • the evaluations of $g^{(q)}$ in the sparse grid points, i.e. the vector $\boldsymbol{g}^{\prime}=(g^{(q)}(\boldsymbol{u}_{1}^{\prime}),\ldots,g^{(q)}(\boldsymbol{u}_{N}^{\prime}))$;
@@
    Create a dataset of pairs $(\boldsymbol{g}^{\prime},\boldsymbol{p})$ using all the vectors collected at the previous item;
@@
    Given a NN with characterizing function $\widehat{\Delta}:\mathbb{R}^{N}\rightarrow[0,1]^{N}$, train the NN with respect to $\mathcal{T}$ and $\mathcal{V}$. In practice, the objective of the training is to obtain a NN that approximates $\Delta$ and, by consequence, the exact discontinuity detector $\Delta^{*}$; i.e.:
@@
    for each evaluation $\boldsymbol{g}^{\prime}$ of a discontinuous function on the grid points of a sparse grid $\mathcal{S}^{\prime}\simeq\mathcal{S}$.
@@
The most expensive part of this procedure is item 3, because it involves a deterministic (approximation of) $\Delta^{*}$; typically, this detector is expensive, especially when we have a combination of a large parameter $t$ and a high space's dimension $n$ (i.e., $t,n\gg 1$).
@@ 4.2 Loss Function
Concerning the loss function for training a NN-based detector, several choices are possible. In [30], the Mean Squared Error is used as loss function for training the convolutional detector. However, since the target vectors $\boldsymbol{p}$ have values in $\{0,1\}$ (not trouble point and troubled point, respectively), we prefer to use the loss function
$\mathcal{L}(\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{(\boldsymbol{g}^{\prime},\boldsymbol{p})}\sum_{i=1}^{N}-w_{1}p_{i}\log(\widehat{p}_{i})-w_{0}(1-p_{i})\log(1-\widehat{p}_{i})\,,$ (19)
for any batch $\mathcal{B}$ of input-output data, and where $\widehat{\boldsymbol{p}}:=\widehat{\Delta}(\boldsymbol{g}^{\prime})\in[0,1]^{N}$.
Specifically, $\mathcal{L}$ is the average sum of the component-wise application of the binary cross-entropy loss. The hyper-parameters $w_{1}$ and $w_{0}$ are used to give more importance to the identification of troubled points or non-troubled points, respectively.
@@ 4.3 NN Models: MLPs and GINNs
Looking at step 6 of the building procedure for NN-based detectors (Section 4.1), we observe that the only practical restrictions for the NN model consists of using an input layer of $N$ units and an output layer of $N$ units with values in $[0,1]$. Therefore, any architecture is allowed for the hidden layers.
@@
Since Algorithm 1 is based on SGGs with same adjacency matrix (see 2.1), it is reasonable to train a NN model tailored for tasks defined on graph-structured data, where the graph is fixed and only the node features vary. In [5], Berrone et al. recently observed that a good choice for learning a target function $\boldsymbol{F}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$, dependent on the adjacency matrix of a fixed graph $G$, is to use a Graph-Informed Neural Network (GINN), because they perform better than Multi-Layer Perceptrons (MLPs) and classic Graph Neural Networks (GNNs). The GINN models are part of the wide family of spatial GNNs and have been specifically developed for such a kind of tasks [5]. Indeed, GNNs are mainly designed for other tasks than learning $\boldsymbol{F}$ (e.g., graph classification, semi-supervised learning on nodes, etc.), while MLPs do not exploit the graph's connections [33, 5].
@@
For these reasons, we focus on MLP and GINN architectures for building the NN-based detectors. However, the numerical experiments in Section 5 confirm the observation of [5]; i.e., we observe that GINN-based detectors perform better than MLP-based detectors.
4.3.1 Graph-Informed Neural Networks
@@
A GINN is a NN characterized by Graph-Informed (GI) Layers [5]. These layers are based on a new variant of the graph-convolution operation. In its simplest version, i.e. in the case of one feature per node, both in input and in output, the action of a GI layer based on a graph with adjacency matrix $A\in\mathbb{R}^{N\times N}$ is a function $\mathcal{L}^{GI}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ such that
$\mathcal{L}^{GI}(\boldsymbol{x})=\boldsymbol{\varphi}\left((A+\mathbb{I}_{N})^{T}\mathrm{diag}(\boldsymbol{w})\,\boldsymbol{x}+\boldsymbol{b}\right)\,,$ (20)
for each vector of input features $\boldsymbol{x}\in\mathbb{R}^{N}$ and where:
  • $\boldsymbol{w}\in\mathbb{R}^{N}$ is the vector containing the $N$ layer weights, associated to the $N$ graph nodes;
@@
  • $\boldsymbol{\varphi}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is the element-wise application of the activation function $\varphi$;
@@
In a nutshell, (20) is equivalent to the action of a ``constrained'' Fully-Connected (FC) layer where the weights are such that
$w_{ij}=\begin{cases}w_{i}\,,\quad&\text{if }a_{ij}\neq 0\text{ or }i=j\\ 0\,,\quad&\text{otherwise}\end{cases}\,.$
Layers characterized by (20) can be generalized to receive any arbitrary number $K\geq 1$ of input features per node and to return any arbitrary number $F\geq 1$ of output features per node. Then, the action of a general GI layer is a function $\mathcal{L}^{GI}:\mathbb{R}^{N\times K}\rightarrow\mathbb{R}^{N\times F}$. Additionally, pooling and mask operations can be added. For more details about GI layers, see [5].
Now, we point the attention of the reader to the fact that the GI layer definition is very general and does not ask for a non-weighted graph. Indeed, one of the main advantages of using a GI layer is that it exploits the zeros of the adjacency matrix $A$, ``pruning'' the connections between layer units according to the graph edges; therefore, the zero elements of $A$ are in the same position independently on the usage of weights for the graph edges.
The observation above is important to build GINNs with respect to the WSGG of a given sparse grid. In particular, we prefer a weighted graph because we want to emphasize the difference between connections according to the distance of the nodes in $\mathbb{R}^{n}$. Then, using a weighted graph for the GINN, we have that the action of the GINN's weights is rescaled with respect to the non-trainable edge weights $\omega_{ij}\in(0,1]$ (see 2.4).
@@ 5 Numerical Experiments
In this section, we apply the proposed NN-based discontinuity detection method (see Algorithm 2) on various test functions, both using MLP-based and GINN-based detectors (for details on architectures, see Section A.4). We recall that our NN-based detectors detect the troubled points reading only the function evaluations at the sparse grid's points; then, they can be applied to any sparse grid (similar to the one used for building the NN model) and to any discontinuous function with domain dimension equal to the one of the sparse grid (see 3.1) or, theoretically, with smaller domain dimension (see 3.5).

### v4

@@ 3.1 Algorithm for Discontinuity Detection
For a fixed $\Lambda_{\min}>0$, let $\kappa\in\mathbb{N},\kappa\geq 0$, be the smallest integer such that $\Lambda/(2^{\kappa+1})<\Lambda_{\min}$; then, $\lambda^{(\kappa)}\leq\frac{\Lambda}{2^{\kappa+1}}<\Lambda_{\min}$ and the algorithm has a finite termination. Moreover, it holds
@@
$\mathrm{dist}(\boldsymbol{x}^{(\kappa)},\delta)\leq\|{\boldsymbol{x}^{(\kappa)}-\boldsymbol{x}^{(\kappa)}_{\delta}}\|\leq\frac{\lambda^{(\kappa)}}{2}<\frac{\Lambda_{\min}}{2}\leq\varepsilon$
and thus the thesis follows.
@@
A discontinuity detector based on a sparse grid $\mathcal{S}$ in $\mathbb{R}^{n}$ can also be used for detecting discontinuities of functions $g^{\prime}:\Omega^{\prime}\subseteq\mathbb{R}^{n^{\prime}}\rightarrow\mathbb{R}$, with $n^{\prime}<n$. Indeed, we can define a function $g:\Omega^{\prime}\times\mathbb{R}^{n-n^{\prime}}\subseteq\mathbb{R}^{n}\rightarrow\mathbb{R}$ as
@@
for each $\boldsymbol{x}\in\Omega^{\prime}\times\mathbb{R}^{n-n^{\prime}}$, and use the detector with respect to $g$. Moreover, in this case Algorithm 1 can be modified to ignore any detected troubled point with at least one coordinate $x_{n^{\prime}+1},\ldots,x_{n}\neq 0$, assuming to start with sparse grids all centered in points with coordinates $x_{n^{\prime}+1}=\cdots=x_{n}=0$. The study of such a kind of generalization is postponed to future work.
@@ 4 NN-based Discontinuity Detectors
First, we will sketch the entire procedure in Section 4.1. Then, in the following subsections, we will provide details about the loss function and NN models. We refer the reader to A for further details related to the building of a deterministic approximation of an exact detector (A.1), the creation of synthetic data for the training (A.2), data preprocessing (A.3), and NN architecture archetypes (A.4).
@@ 4.1 Building Procedure for NN-based Detectors
    Choose a sparse grid $\mathcal{S}$ in $\mathbb{R}^{n}$. Let $N$ denote the number of points in $\mathcal{S}$.
@@
    For each $g^{(q)}\in\mathcal{G}$, run Algorithm 1 with either an exact detector $\Delta^{*}$ or a deterministic approximation $\Delta$ of $\Delta^{*}$ (see A.1 for details); then, at each detection step with respect to a sparse grid $\mathcal{S}^{\prime}\simeq\mathcal{S}$, compute:
    • the evaluations of $g^{(q)}$ in the sparse grid points, i.e. the vector $\boldsymbol{g}^{\prime}=(g^{(q)}(\boldsymbol{x}_{1}^{\prime}),\ldots,g^{(q)}(\boldsymbol{x}_{N}^{\prime}))$;
@@
    Create a dataset of pairs $(\boldsymbol{g}^{\prime},\boldsymbol{p})$ collecting the vectors obtained at the previous step;
@@
    Given a NN with characterizing function $\widehat{\Delta}:\mathbb{R}^{N}\rightarrow[0,1]^{N}$, train the model with respect to $\mathcal{T}$ and $\mathcal{V}$ to obtain an approximation of $\Delta$ and, as a consequence, the exact discontinuity detector $\Delta^{*}$; i.e.:
@@
    for each vector $\boldsymbol{g}^{\prime}$ of function evaluations at the grid points of $\mathcal{S}^{\prime}$.
@@
The most expensive part of this procedure is item 3, because it involves a deterministic (approximation of) $\Delta^{*}$; typically, this detector is expensive, especially when we have a high space’s dimension $n$ (i.e., $n\gg 1$).
@@ 4.2 Loss Function
Concerning the loss function for training a NN-based detector, several choices are possible. In [3], the Mean Squared Error is used as loss function for training the convolutional detector. However, since the target vectors $\boldsymbol{p}$ have values in $\{0,1\}$ (non-trouble point and troubled point, respectively), we prefer to use the loss function
$\mathcal{L}(\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{(\boldsymbol{g}^{\prime},\boldsymbol{p})}\sum_{i=1}^{N}-{\mu_{1}}p_{i}\log(\widehat{p}_{i})-{\mu_{0}}(1-p_{i})\log(1-\widehat{p}_{i})\,,$ (17)
for any batch $\mathcal{B}$ of input-output data, and where $\widehat{\boldsymbol{p}}:=\widehat{\Delta}(\boldsymbol{g}^{\prime})\in[0,1]^{N}$; namely, $\mathcal{L}$ is the average sum of the component-wise application of the binary cross-entropy loss [32, 33]. The hyper-parameters $\mu_{1}$ and $\mu_{0}$ are used to suitably weight the identification of troubled points or non-troubled points, respectively.
@@ 4.3 NN Models: MLPs and GINNs
Looking at step 6 of the building procedure for NN-based detectors (Section 4.1), we observe that the only practical restrictions for the NN model consist of using an input layer of $N$ units and an output layer of $N$ units with values in $[0,1]$. Therefore, any architecture is allowed for the hidden layers.
@@
Since Algorithm 1 is based on SGGs with the same adjacency matrix (see Proposition 2.1), it is reasonable to train a NN model tailored for tasks defined on graph-structured data, where the graph is fixed and only the node features vary. In [26], Berrone et al. recently observed that a good choice for learning a target function $\boldsymbol{F}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$, dependent on the adjacency matrix of a fixed graph $G$, is to use a Graph-Instructed Neural Network (GINN) because they perform better than Multi-Layer Perceptrons (MLPs) and classic Graph Neural Networks (GNNs). The GINN models are part of the wide family of spatial GNNs and have been specifically developed for such a kind of task [26]. Indeed, GNNs are mainly designed for other tasks than learning $\boldsymbol{F}$ (e.g., graph classification, semi-supervised learning on nodes, etc.), while MLPs do not exploit the graph’s connections [34, 26, 27]. We point the reader to the fact that in [26] GI layers and GINNs are denoted as Graph-Informed layers and Graph-Informed NNs, respectively. In [35], in a different framework from the one addressed in [26], a homonymous but different model is presented; therefore, starting from the work [27] of Della Santa et al., the authors changed the names of both layers and NNs, to avoid confusion with [35].
@@
Due to the observations about MLPs and GINNs, we focus here on these architectures for building the NN-based detectors. Moreover, the numerical experiments in Section 5 confirm the observation of [26]; i.e., we observe that GINN-based detectors perform better than MLP-based detectors.
4.3.1 Graph-Instructed Neural Networks
A GINN is a NN characterized by Graph-Instructed (GI) Layers. These layers are defined by an alternative graph-convolution operation introduced in [26]. In brief, from a message-passing point of view, this graph-convolution operation is equivalent to having each node $v_{i}$ of $G$ sending to its neighbors a message equal to the input feature $x_{i}$, scaled by a weight $w_{i}$; then, each node sum up all the messages received from the neighbors, add the bias, and applies the activation function. We point the attention of the reader to the fact that, only in this subsection, $\boldsymbol{x}$ denotes feature vectors and not the sparse grid’s points.
In a nutshell, the message-passing interpretation can be summarized by Figure 5 and the following node-wise equation
$x_{i}^{\prime}=\sum_{j\in\mathrm{N}_{\text{in}}(i)\cup\{i\}}x_{j}\,w_{j}+b_{i}\,,$ (18)
where $x_{i}^{\prime}$ is the layer’s output feature corresponding to node $v_{i}$ and $\mathrm{N}_{\rm in}(i)$ is the set of indices such that $j\in\mathrm{N}_{\rm in}(i)$ if and only if $e_{ij}=\{v_{i},v_{j}\}$ is an edge of the graph. We dropped the action of the activation function for simplicity.
[…]
Figure 5: Visual representation of (18). Example with $n=4$ nodes (non-directed graph), $i=1$; for simplicity, the bias is not illustrated.
@@ 4.3.1 Graph-Instructed Neural Networks
Starting from (18), we can formally define the GI layers through a compact formulation. Given a graph $G$ (without self-loops) and its adjacency matrix $A\in\mathbb{R}^{N\times N}$, the simplest version of GI layer for $G$ is a NN layer described by a function $\mathcal{L}^{GI}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$, with one input feature per node and one output feature per node, such that
[…]
where $\boldsymbol{x}\in\mathbb{R}^{N}$ denotes the vector of input features and:
  • $\boldsymbol{w}\in\mathbb{R}^{N}$ is the weight vector, with the component $w_{i}$ associated to the graph node $v_{i}$, $i=1,\ldots,N$;
  • $\mathrm{diag}(\boldsymbol{w})\in\mathbb{R}^{N\times N}$ is the diagonal matrix with elements of $\boldsymbol{w}$ on the diagonal and $\mathbb{I}_{N}\in\mathbb{R}^{N\times N}$ is the identity matrix. For future reference, we set $\widehat{W}:=\mathrm{diag}(\boldsymbol{w})(A+\mathbb{I}_{N})$; the diagonal matrix $\mathrm{diag}(\boldsymbol{w})$ in (19) is introduced only for describing in matrix form the multiplication of the $i$-th row of $(A+\mathbb{I}_{N})$ by the weight $w_{i}$ associated to node $v_{i}$, for each $i=1,\ldots,N$;
@@
  • $\boldsymbol{\psi}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is the element-wise application of the activation function $\psi$;
@@
From another point of view, equation (19) is equivalent to the action of a “constrained” Fully-Connected (FC) layer where the weights are the same if the connection is outgoing from the same unit, whereas it is zero if two units correspond to graph nodes that are not connected (see Figure 6); more precisely:
$\widehat{w}_{ij}=\begin{cases}w_{i}\,,\quad&\text{if }a_{ij}\neq 0\text{ or }i=j\\ 0\,,\quad&\text{otherwise}\end{cases}\,,$ (20)
where $a_{ij},\widehat{w}_{ij}$ denote the $(i,j)$-th element of $A,\widehat{W}$, respectively.
[…]
(a) FC representation of GI layer
$\widehat{W}^{T}=\begin{bmatrix}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}w_{1}}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}w_{2}}&0&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}w_{4}}\\ {\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}w_{1}}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}w_{2}}&{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}w_{3}}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}w_{4}}\\ 0&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}w_{2}}&{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}w_{3}}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}w_{4}}\\ {\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}w_{1}}&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}w_{2}}&{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}w_{3}}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}w_{4}}\end{bmatrix}$
(b) Weight matrix of GI layer
Figure 6: Visual representation of a GI layer as a “constrained” FC layer (subfigure ($a$)), with weight matrix defined by (20) (subfigure ($b$)). This figure is based on the same graph illustrated in Figure 5
Layers characterized by (19) can be generalized to read any arbitrary number $K\geq 1$ of input features per node and to return any arbitrary number $F\geq 1$ of output features per node. Then, the action of a general GI layer is a function $\mathcal{L}^{GI}:\mathbb{R}^{N\times K}\rightarrow\mathbb{R}^{N\times F}$. Additionally, pooling and mask operations can be added. For more details about GI layers, see [26, 27, 36].
Now, we point the attention of the reader to the fact that the GI layer definition is very general and does not ask for a non-weighted graph. Indeed, one of the main advantages of using a GI layer is that it exploits the zeros of the adjacency matrix $A$, “pruning” the connections between layer units according to the graph edges; therefore, the zero elements of $A$ are in the same position independently on the usage of weights for the graph edges.
The observation above is important to build GINNs with respect to the WSGG of a given sparse grid. In particular, we prefer a weighted graph because we want to emphasize the difference between connections according to the distance of the nodes in $\mathbb{R}^{n}$. Then, using a weighted graph for the GINN, we have that the action of the GINN’s weights is rescaled with respect to the non-trainable edge weights $\omega_{ij}\in(0,1]$ (see Definition 2.4).
@@ 5 Numerical Experiments
In this section, we apply the proposed NN-based discontinuity detection method (see Algorithm 2) on various test functions, both using MLP-based and GINN-based detectors (for details on architectures, see A.4). We recall that our NN-based detectors detect the troubled points reading only the function evaluations at the sparse grid’s points; then, they can be applied to any sparse grid (similar to the one used for building the NN model) and to any discontinuous function with domain dimension equal to the one of the sparse grid (see Remark 3.1) or, in principle, with smaller domain dimension (see Remark 3.5).